OpenStudy (jojo12g):

During a period of 6 months, MegaCorp had monthly profits of $3.2 million, $1.8 million, and $6.4 million. Two months had losses of $4.5 million and $2.7 million. The net profit was $7.5 million. What happened in the month that is not included in the problem data? A. The company had a loss of $3.3 million. B. The company had a profit of $3.3 million. C. The company had a loss of $11.7 million. D. The company had a profit of $11.7 million.

OpenStudy (jojo12g):

please help give ribbon

OpenStudy (envy_seven):

The answer is B

OpenStudy (jojo12g):

thax

OpenStudy (envy_seven):

3.2 + 1.8 + 6.4 = 11.4 11.4 - (4.5 + 2.7) = 4.2 4.2 + 3.3 = 7.5 :)

OpenStudy (envy_seven):

Thanks for the medal

OpenStudy (jojo12g):

welcome